Circular No. 7889

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J0929–314 J. G. Greenhill, A. B. Giles, and K. M. Hill, University of Tasmania,

report a possible optical counterpart for XTE J0929–314 (cf. IAUC 7888). Observations obtained around May 1.42–1.58 UT at the 1-m Mt. Canopus telescope show a blue object with V ∼ 18.8 at α = 9^{h}29^{m}20^{s}.16, δ = −31^{o}23′02′′.7 (equinox J2000.0; uncertainty ± 0′′.5). The object was also detected in B, R, and I, but not on a red plate from the Digitized Sky Survey. The source faded significantly during the observations.

V838 MONOCEROTIS A. Henden, Universities Space Research Association and U.S. Naval

Observatory (USNO); U. Munari, P. Marrese, and F. Boschi, Osservatorio Astrofisico di Asiago; and R. Corradi, Isaac Newton Group of Telescopes, La Palma, report that the light echo surrounding V838 Mon is now developing an inner edge on U-band images obtained with the USNO 1-m telescope, ∼ 6′′ radius from the star on Apr. 30.2 UT and propogating outwards. This is caused by the large and fast drop in U-band brightness at 0.17 mag/day that V838 Mon has experienced over the last few days: Apr. 18.12, U = 14.09; 22.14, 14.76; 30.13, 16.09. Continuing the trend, the light echo should disappear by late May, when the fading front will have reached the outer 34′′ radius that has held firm for the past month. High-resolution échelle CCD spectra obtained with the Asiago 1.82-m telescope on Apr. 22.81 show the development of M-giant absorption features. Narrow P-Cyg profiles still persist for several metallic lines and Na I D, while the Ca II near-infrared triplet is now in pure absorption. The only significant emission feature is Hα with a multipeaked complex profile of 220 km/s half-width and −5.5 km/s heliocentric radial velocity for the strongest peak. Remarkable TiO bands in absorption resemble an M6/7 spectrum. No carbon-related absorption bands are evident.

Visual magnitude estimates: Mar. 5.79 UT, 7.4 (K. Sárneczky, Piszkéstető, Hungary); 8.441, 7.0 (S. Yoshida, Ibaraki, Japan); 11.860, 7.7 (M. Martignoni, Busto Arsizio, Italy); 15.82, 7.4 (A. R. Baransky, Kiev, Ukraine); 19.067, 7.5 (R. Y. Shida, São Paulo, Brazil); 26.793, 8.0 (K. Hornoch, Lelekovice, Czech Republic); Apr. 2.800, 8.2 (Hornoch); 4.950, 8.0 (Shida); 10.038, 8.4 (Shida); 12.034, 8.7 (Shida); 15.968, 9.0 (Shida).

2002 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green